OpenStudy (bojack97):

Help please, I will give medal and fan. The sets of numbers 7, 24, 25 and 9, 40, 41 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples.

OpenStudy (yshreya50):

Pythegorean triplets means sum of squares of two smallest number equals square of largest number amongst three. Much like in pythegorus theorem 7^2=49 24^2=576 25^2=625 now 49+576 = 625. This helps in geometry calculation if you know value of two side you won't have to calculate much to find third side.

OpenStudy (bojack97):

Im still a little confused...

OpenStudy (wolf1728):

A Pythagorean Triple means three numbers a, b and c such that the sum of a^2 + b^2 = c2 In order to be a Pythagorean Triple, the numbers have to be integers. For example Pythagorean Triples are: 3, 4 and 5 and 6, 8 and 10 5, 12 and 13 6, 7 and 9.2195444573 are NOT Pythagorean Triples because 9.2195444573 is NOT an integer

OpenStudy (bojack97):

So any integer is a triple?

OpenStudy (wolf1728):

No, not any integer is NOT a triple. When you have 3 numbers a, b and c that have a^2 + b^2 = c^2 it is a Pythagorean Triple if all 3 numbers are integers

OpenStudy (bojack97):

So the answer to my question, The sets of numbers 7, 24, 25 and 9, 40, 41 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. is because both sets are all integers?

OpenStudy (bojack97):

or all 3 numbers in both sets are integers...

OpenStudy (wolf1728):

7, 24 and 25 and 9, 40 and 41 are BOTH Pythagorean Triples because 1) a^2 + b^2 = c^2 AND 2) all three numbers are integers

OpenStudy (wolf1728):

You need BOTH condition 1 and 2 to be trues for Pythagorean Triples

OpenStudy (wolf1728):

You just can't pick any 3 integers, let's say 17, 18 and 19 They are all integers but 17^2 + 18^2 does NOT equal 19^2

OpenStudy (bojack97):

Alright Thank you so much!

OpenStudy (wolf1728):

u r welcome :-)